o-eA uniform electric field is shown below. Draw equipotential surfaces. In which direction does the voltage decrease? If an electron is placed in the electric field at rest, draw the direction of the force. The electron moves from higher potential point to lower potential point or the other way around? If the potential difference between two locations of the electron is 5V, what is the change in potential energy of the electron. What is its speed at the end of the motion through 5V?
Answer #1

Voltage decreases in the direction of the E field.

The direction of the force on an electron is opposite to the direction of E field.

Electron moves from higher potential point to lower potential point.

Change in PE = e*dV

= (1.6*10^-19)(5)

= 8*10^-19 V

KE = PE

=> 1/2*mv^2 = 8*10^-19

=> 1/2*9.11*10^-31*v^2 = 8*10^-19

=> v = 1.33*10^6 m/s